OpenStudy (anonymous):

Tickets are numbered from 1 to 10. Two tickets are drawn one after the other at random. Find the probability that the number on one of the tickets is a multiple of 5 and on the other a multiple of 4.

OpenStudy (princeharryyy):

multiples of 5,10 total tickets =10, possibilities for multiple of 5 =2 so, 2/10 = 1/5

OpenStudy (princeharryyy):

for 4 sample space left =9, possibilities 4,8 i.e. 2 so, ans =2/9

OpenStudy (princeharryyy):

r u getting it @cutelilgirl